# An empirical study on code review activity prediction in practice

### v1

@@ Abstract.
ccs: Software and its engineering Software maintenance tools
@@ 1. Introduction
In software development, code reviewing is an essential quality assurance step during code integration (rigby2012contemporary, ; fagan2002history, ). Developers submit their code for review as a patch. Then, reviewers either validate the code change, which is then forwarded to integration, or they provide feedback. The review process can go through several iterations, where the reviewer comments on the submission to have the developer revise it and submit a new version, referred to as a revision. Eventually, if the developer addresses all comments, the patch is accepted, otherwise it is rejected.
Given the essential role of code reviews in modern development (sadowski2018modern, ), it is crucial to have a high code review velocity and effectiveness. Yet, Hong et al. report that developers waited a median of 15-64 hours for reviewers’ feedback in three open source projects (i.e., Openstack Nova, Openstack Ironic and Qt Base) (hong2022should, ). Those numbers were confirmed by our industrial partner company, which has similar waiting times. Such long waiting times delay the integration of code and thus the projects’ development and release time. Furthermore, other factors can cause review processes to take a long time, including the reviewers involved (e.g., availability and experience), the submitted code (e.g., patch size and complexity, developers’ experience) and the number of review iterations needed (kononenko2016code, ). Reviewers are usually developers on the project, who have to split their work time between their own development tasks and reviews. Furthermore, they might have to review many patches, themselves composed of many files and changes.
In order to improve review effectiveness, Hong et al. (hong2022should, ) automate the identification of the parts of patches that either need revisions or comments using a Bag-of-Words (BoW) embedding and two review process features (i.e., number of lines added and removed). As BoW embeddings do not represent the order of words such as Large Language Model (LLM) embeddings preferred in recent NLP-based tasks (hou2023large, ) do, we suspect that the former do not capture enough context. With approaches such as risk/bug prediction using software process features (nayrolles2018clever, ; giger2012method, ; kamei2012large, ), it is also unknown how that type of embedding performs in the context of review activities.
@@
Prediction of review activity needs has potential to assist reviewers through multiple use cases, such as adding warnings in the code for reviewers to focus on (henley2018cfar, ) or re-ordering files (huang2018salient, ; petrovic2018state, ). In this paper, we focused on the latter. In a wizard-of-oz user study with 29 professional developers, we find that, inspired by Fregnan et al.’s suggestion to “display first those files that are more critical” (fregnan2022first, ), hot-spot based file re-ordering correlates with improved review comment quality. In particular, more review comments were provided (+23%), and those comments were better targeted towards known hot-spots (+13% precision and +8% recall) compared to the alphanumeric ordering. Even though reviewers would still be required to look at the rest of a patch, automated identification of hot-spots would attract the reviewer’s attention earlier to the most critical parts.
@@
  • A replication package with the open-source dataset (replic, ).
@@ 2. Background and Related Work
For years, the software engineering domain has proposed approaches to help developers in tasks such as defect prediction (borg2019szz, ; nayrolles2018clever, ), code recommendation (takuya2011spontaneous, ), and others (kim2013should, ; xia2017effective, ; da2016framework, ; anvik2006should, ). In the last 10 years, the focus has moved towards helping reviewers as well, regarding subjects such as reviewer recommendation (thongtanunam2015should, ; rahman2016correct, ; thongtanunam2014improving, ), code change ordering (baum2017optimal, ), code diff quality (li2022automating, ; hong2022should, ), review comment recommendation (hong2022commentfinder, ; li2022automating, ; tufano2022using, ; siow2020core, ; gupta2018intelligent, ; tufano2021towards, ), review comment evaluation (bosu2015characteristics, ; rahman2017predicting, ; efstathiou2018code, ) and code refinement (li2022automating, ; tufano2021towards, ; tufano2022using, ). In this paper, we focus on the subject of code diff quality, by identifying code submissions needing review activity (i.e., revisions and/or comments). Bosu et al. reported that 34.50% of code review comments from five major Microsoft projects that they studied are not “useful”, i.e., did not lead to code modification, thus arguing the need for code review evaluation (bosu2015characteristics, ). However, we consider that any review activity has a high value. Submission of a revision indicates that the code needed to be modified due to an issue (e.g., bug or efficiency) or did not follow some code convention of the team. Review comments have also value as they might indicate unclarity in the code (e.g., the nomenclature is not clear), the need for changes in another part of the submission, or the need for a future code submission treated separately. As both revisions and comment have value, we also want to predict ”hot-spots” as changes that requires any of those review activities.
@@
Li et al. explored code diff quality in order to identify parts of the code change that need review comments (li2022automating, ). Their approach uses a Text-To-Text Transfer Transformer (T5) that takes as input the code change and returns a probability that the code change needs a comment. Their T5 is pre-trained on a dataset composed of on code diffs. Hong et al. proposed the “RevSpot” approach to predict problematic lines, i.e., lines that will be commented on by the reviewer or revised (hong2022should, ). Their approach uses input based on code vocabulary and they train classification models such as Random Forest and Naive Bayes. This simpler training approach has the advantage of being less costly, whereas fine-tuning a T5 model is known to be computationally expensive. The classification-based approach reports similar results on the ”commented” prediction task, indicating that the computational effort of fine-tuning an LLM might not be relevant for this task.
@@
We also explore in this paper the use of Large Language Model (LLM) encoders for text embedding. LLM are models trained to solve Natural Language tasks. They distinguish themselves from Language Models (LM) due to the size large number of parameters if the LLM models (hou2023large, ). The need for large models comes from the observation that the more parameters are added, the better the LM model performs (wei2022emergent, ). LLMs are used in three main types of architectures: encoder-only, encoder-decoder and decoder-only. Encoder-only architectures take text as input and generate an embedding, i.e., an abstracted representation of the code in the form of a vector capturing the text’s semantics (devlin2018bert, ). Encoder-decoder architectures take text as input, then generate an intermediate embedding state that is itself fed as input to the decoder (vaswani2017attention, ; han2021transformer, ). The intermediate state is used as a way to represent the connection between various inputs and outputs. This type of architecture is usually used for translation tasks or question-answering. Finally, the decoder-only architectures use a decoder to generate text, usually in a sequential prediction paradigm. This type of model has been vastly popularized over the last few years with GPT models (openai2023gpt4, ) and ChatGPT (chatgpt, ).
@@
LLMs have been used in the context of Software Engineering for multiple tasks, such as code synthesis/generation (austin2021program, ; zan2023large, ), review automation (tufano2021towards, ; tufano2022using, ; li2022automating, ), and code fuzzing (deng2022fuzzing, ). In this paper, we focus on the encoder-only type of task, as we use the LLM as a text embedding approach. The advantage of this is that we can use a pre-trained LLM model as a zero-shot encoder, where we do not fine-tune the LLM. The embedding is then used as input to a classifier, easier to train, thus giving a more cost-effective and lower computation-time training approach. Previous studies have found ”simpler” approaches such as classification models to be similar, and sometimes even better, than Deep Learning (DL) approaches in the context of software engineering tasks in terms of both performance (fu2017easy, ; hellendoorn2017deep, ) and speed (majumder2018500, ), which motivates our use of classifiers on top of a zero-shot LLM embeddings instead of fine-tuning a LLM encoder-decoder model.
@@
Other past works regarding bug prediction (giger2012method, ), risk prediction (shihab2012industrial, ; nayrolles2018clever, ) and quality assurance (kamei2012large, ) have shown good performance by using software process features. Those features range from size-based features (e.g., number of lines added, removed), contextual features (e.g., size of the patch), and history-based features (e.g., number of times a file was changed, number of participants on a file), to experience-based features (e.g., experience of developers/reviewers). However, review process prediction tasks have not considered exploring those types of features yet, apart from Hong et al. considering only the two following features: number of lines added and the number of lines removed (hong2022should, ). Exploring review process features approaches is thus one of the goals of this paper.
@@ 3. Motivational Use Case
By default, most review tools list the changed files of a patch in alphanumeric order, even though previous work argues that this might be suboptimal, especially for long code submissions (baum2016need, ; barnett2015helping, ). Yet, it is common, especially in closed-source datasets, to have many file changes in a patch, making the review process more difficult and time consuming. Case in point, across the five datasets studied in section 5, we observed that 38-68% of patches are composed of more than 2 files, while in the two closed-source datasets, 25-30% of patches have more than 5 files.
Baum et al. established principles for code change ordering to help reviewers become more effective (baum2017optimal, ), while Li et al. (li2022automating, ) and Hong et al. (hong2022should, ) predict submission (patch) parts that need either review comments or a revision, yet neither of them explore or discuss the use of code diff quality as a way to prioritize the code change order. Fregnan et al.’s user study indicated that the earlier a file is shown in a patch, the more likely it is to be commented on (fregnan2022first, ), hence suggesting to prioritize files that need review activity to the beginning of a patch.
@@ 3.1. User study set-up
We carried out a wizard-of-oz user study (dahlback1993wizard, ) to compare hot-spot-based ordering of a patch’s changed files to the typical alphanumeric ordering  [^1] The study was approved by an ethical board. The certificate number is not disclosed to respect double-blind.. The wizard-of-oz design means that we used knowledge about the actual hot-spots in the oracle (obtained similarly as in Section 4.1), which is useful to evaluate the promise of hot-spot prioritization for file reordering in optimal conditions compared to the alphanumeric ordering used by most review tools (fregnan2022first, ). The user study was held on the premises of our industrial partner, and involved 29 developers with an average experience at the company of 6 years, knowledgeable both in developing and reviewing for the studied closed-source project, which is actively developed by more than 800 developers, with 3-4k patches per month.
@@
Each set was then associated with two patches (from the participants’ common subsystems) that none of the participants had submitted or reviewed before. We selected only patches of exactly five files, similar to Fregnan et al. (fregnan2022first, ). This choice is motivated by the need for patches large enough to observe the impact of ordering files while keeping the review effort manageable to attract more participants. The selected patches were not older than 2 months prior to the study, had received reviews, and needed at least one later revision or involved at least one thread of review comments.
@@ 4.1. Labeling
Our review activity prediction tasks extend existing code diff quality prediction works (li2022automating, ; hong2022should, ) in order to highlight changed files that most likely should be commented on or revised. Past approaches have used a binary labeling based on if the changed file was commented on or revised (li2022automating, ; hong2022should, ). Some works have also addressed the question of identifying useful comments, as they consider that commented changes that do not lead to revisions are not useful (bosu2015characteristics, ). However, we consider such comments to still have value, as they can lead to future code submission or discussions.
Hence, in this paper, we consider similar labels (”commented” and ”revised”) as prior works, but also define a ”hot-spot” label, which indicates whether a code change has any review activity (i.e., comment, revisions, or both). This is because, while comments are meant to lead to revisions, and revisions to be caused by comments, review comments might indicate the need for revisions elsewhere in the patch or changes in future independent patches. Furthermore, we observe that 10-17% of patches across our datasets contain changes that are revised, while never commented, typically because of a review comment made on another file or outside the reviewing platform. If considering only patches that were commented on, those cases would thus be missed, hence the importance of considering hot-spot prediction. The labels’ definitions can be found in Table 3. This labeling is done at a file-level granularity, following Hong et al.’s approach (hong2022should, ).
@@ 4.2. Datasets
In this paper, we use three open source datasets and two closed source datasets provided by our industrial partner (different from the dataset used in the user study). The open source datasets we consider are those of Hong et al. (hong2022should, ), i.e., Qt Base (Base), Openstack Ironic (Ironic), Openstack Nova (Nova), which were used for commented and revisioned prediction tasks (hong2022should, ; tufano2021towards, ). We then used Gerrit’s API to gather extra information about the changed files, i.e., author id, reviewer ids, and full file diff (since their replication package only has the added lines), needed for our embeddings.
@@ 4.3. Data splitting
In the case of the open-source datasets, we split each dataset along its timeline into a training set validation set and test set of size 80%-10%-10% respectively, similarly to Hong et al.’s approach (hong2022should, ).
@@ 4.4.1. Bag-of-Words
The BoW approach represents each input as a vector of word frequency where each column represents a word in the corpus vocabulary and the value in the vector at that column indicates the number of occurrences of the word in the input. Considering code changes as input, we have to clean the input in this context, since the source code of changes have many symbols that are not text. We follow the state-of-the-art’s approach (hong2022should, ) for code pre-processing, which only considers alphanumeric terms. Numbers are also replaced by a token ”<NUMBER>”. Still following this approach, we do not perform any stemming, lowercase or lemmatization transformation as such approaches are valuable for natural languages used by humans but do not make sense in the case of code. Regarding the BoW computation itself, we use the Countvectorize function of the Scikit-Learn Python library (contvect, ). The Countvectorize function is trained on the training set, then is applied to the test set for each data split. Words that appear in the test set but not in the training set will thus not be included in the bag of words. To reduce the vocabulary size, we remove words that appear only in one single changed files in the corpus as well as words that are too common (i.e., appear in more than 50% of the changed file corpus).

### v2

# An Empirical Study on Code Review Activity Prediction and Its Impact in Practice

@@ Abstract.
copyright: rightsretaineddoi: 10.1145/3660806journalyear: 2024submissionid: fse24main-p790-pjournal: PACMSEjournalvolume: 1journalnumber: FSEarticle: 99publicationmonth: 7ccs: Software and its engineering Software maintenance tools
@@ 1. Introduction
In software development, code reviewing is an essential quality assurance step during code integration (Rigby et al., 2012; Fagan, 2002). Developers submit their code for review as a patch. Then, reviewers either validate the code change, which is then forwarded to integration, or they provide feedback. The review process can go through several iterations, where the reviewer comments on the submission to have the developer revise it and submit a new version, referred to as a revision. Eventually, if the developer addresses all comments, the patch is accepted, otherwise it is rejected.
Given the essential role of code reviews in modern development (Sadowski et al., 2018), it is crucial to have a high code review velocity and effectiveness. Yet, Hong et al. report that developers waited a median of 15-64 hours for reviewers’ feedback in three open source projects (i.e., Openstack Nova, Openstack Ironic and Qt Base) (Hong et al., 2022a). Those numbers were confirmed by our industrial partner company, which has similar waiting times. Such long waiting times delay the integration of code and thus the projects’ development and release time. Furthermore, other factors can cause review processes to take a long time, including the reviewers involved (e.g., availability and experience), the submitted code (e.g., patch size and complexity, developers’ experience) and the number of review iterations needed (Kononenko et al., 2016). Reviewers are usually developers on the project, who have to split their work time between their own development tasks and reviews. Furthermore, they might have to review many patches, themselves composed of many files and changes.
In order to improve review effectiveness, Hong et al. (Hong et al., 2022a) automate the identification of the parts of patches that either need revisions or comments using a Bag-of-Words (BoW) embedding and two review process features (i.e., number of lines added and removed). As BoW embeddings do not represent the order of words such as Large Language Model (LLM) embeddings preferred in recent NLP-based tasks (Hou et al., 2023) do, we suspect that the former do not capture enough context. With approaches such as risk/bug prediction using software process features (Nayrolles and Hamou-Lhadj, 2018; Giger et al., 2012; Kamei et al., 2012), it is also unknown how that type of embedding performs in the context of review activities.
@@
Prediction of review activity needs has potential to assist reviewers through multiple use cases, such as adding warnings in the code for reviewers to focus on (Henley et al., 2018) or re-ordering files (Huang et al., 2018; Petrović and Ivanković, 2018). In this paper, we focused on the latter. In a wizard-of-oz user study with 29 professional developers, we find that, inspired by Fregnan et al.’s suggestion to “display first those files that are more critical” (Fregnan et al., 2022), hot-spot based file re-ordering correlates with improved review comment quality. In particular, more review comments were provided (+23%), and those comments were better targeted towards known hot-spots (+13% precision and +8% recall) compared to the alphanumeric ordering. Even though reviewers would still be required to look at the rest of a patch, automated identification of hot-spots would attract the reviewer’s attention earlier to the most critical parts.
@@
  • A replication package with the open-source dataset (rep, [n. d.]).
@@ 2. Background and Related Work
For years, the software engineering domain has proposed approaches to help developers in tasks such as defect prediction (Borg et al., 2019; Nayrolles and Hamou-Lhadj, 2018; Olewicki et al., 2024), code recommendation (Takuya and Masuhara, 2011), and others (Kim et al., 2013; Xia and Lo, 2017; Da Costa et al., 2016; Anvik et al., 2006; Olewicki et al., 2022). In the last 10 years, the focus has moved towards helping reviewers as well, regarding subjects such as reviewer recommendation (Thongtanunam et al., 2015; Rahman et al., 2016; Thongtanunam et al., 2014), code change ordering (Baum et al., 2017), code diff quality (Li et al., 2022; Hong et al., 2022a), review comment recommendation (Hong et al., 2022b; Li et al., 2022; Tufano et al., 2022; Siow et al., 2020; Gupta and Sundaresan, 2018; Tufano et al., 2021), review comment evaluation (Bosu et al., 2015; rahman2017predicting; Efstathiou and Spinellis, 2018) and code refinement (Li et al., 2022; Tufano et al., 2021, 2022). In this paper, we focus on the subject of code diff quality, by identifying code submissions needing review activity (i.e., revisions and/or comments). Bosu et al. reported that 34.50% of code review comments from five major Microsoft projects that they studied are not “useful”, i.e., did not lead to code modification, thus arguing the need for code review evaluation (Bosu et al., 2015). However, we consider that any review activity has a high value. Submission of a revision indicates that the code needed to be modified due to an issue (e.g., bug or efficiency) or did not follow some code convention of the team. Review comments have also value as they might indicate unclarity in the code (e.g., the nomenclature is not clear), the need for changes in another part of the submission, or the need for a future code submission treated separately. As both revisions and comment have value, we also want to predict ”hot-spots” as changes that requires any of those review activities.
@@
Li et al. explored code diff quality in order to identify parts of the code change that need review comments (Li et al., 2022). Their approach uses a Text-To-Text Transfer Transformer (T5) that takes as input the code change and returns a probability that the code change needs a comment. Their T5 is pre-trained on a dataset composed of on code diffs. Hong et al. proposed the “RevSpot” approach to predict problematic lines, i.e., lines that will be commented on by the reviewer or revised (Hong et al., 2022a). Their approach uses input based on code vocabulary and they train classification models such as Random Forest and Naive Bayes. This simpler training approach has the advantage of being less costly, whereas fine-tuning a T5 model is known to be computationally expensive. The classification-based approach reports similar results on the ”commented” prediction task, indicating that the computational effort of fine-tuning an LLM might not be relevant for this task.
@@
We also explore in this paper the use of Large Language Model (LLM) encoders for text embedding. LLM are models trained to solve Natural Language tasks. They distinguish themselves from Language Models (LM) due to the size large number of parameters if the LLM models (Hou et al., 2023). The need for large models comes from the observation that the more parameters are added, the better the LM model performs (Wei et al., 2022). LLMs are used in three main types of architectures: encoder-only, encoder-decoder and decoder-only. Encoder-only architectures take text as input and generate an embedding, i.e., an abstracted representation of the code in the form of a vector capturing the text’s semantics (Devlin et al., 2018). Encoder-decoder architectures take text as input, then generate an intermediate embedding state that is itself fed as input to the decoder (Vaswani et al., 2017; Han et al., 2021). The intermediate state is used as a way to represent the connection between various inputs and outputs. This type of architecture is usually used for translation tasks or question-answering. Finally, the decoder-only architectures use a decoder to generate text, usually in a sequential prediction paradigm. This type of model has been vastly popularized over the last few years with GPT models (OpenAI, 2023) and ChatGPT (cha, [n. d.]).
@@
LLMs have been used in the context of Software Engineering for multiple tasks, such as code synthesis/generation (Austin et al., 2021; Zan et al., 2023), review automation (Tufano et al., 2021, 2022; Li et al., 2022), and code fuzzing (Deng et al., 2022). In this paper, we focus on the encoder-only type of task, as we use the LLM as a text embedding approach. The advantage of this is that we can use a pre-trained LLM model as a zero-shot encoder, where we do not fine-tune the LLM. The embedding is then used as input to a classifier, easier to train, thus giving a more cost-effective and lower computation-time training approach. Previous studies have found ”simpler” approaches such as classification models to be similar, and sometimes even better, than Deep Learning (DL) approaches in the context of software engineering tasks in terms of both performance (Fu and Menzies, 2017; Hellendoorn and Devanbu, 2017) and speed (Majumder et al., 2018), which motivates our use of classifiers on top of a zero-shot LLM embeddings instead of fine-tuning a LLM encoder-decoder model.
@@
Other past works regarding bug prediction (Giger et al., 2012), risk prediction (Shihab et al., 2012; Nayrolles and Hamou-Lhadj, 2018) and quality assurance (Kamei et al., 2012) have shown good performance by using software process features. Those features range from size-based features (e.g., number of lines added, removed), contextual features (e.g., size of the patch), and history-based features (e.g., number of times a file was changed, number of participants on a file), to experience-based features (e.g., experience of developers/reviewers). However, review process prediction tasks have not considered exploring those types of features yet, apart from Hong et al. considering only the two following features: number of lines added and the number of lines removed (Hong et al., 2022a). Exploring review process features approaches is thus one of the goals of this paper.
@@ 3. Motivational Use Case
By default, most review tools list the changed files of a patch in alphanumeric order, even though previous work argues that this might be suboptimal, especially for long code submissions (Baum and Schneider, 2016; Barnett et al., 2015). Yet, it is common, especially in closed-source datasets, to have many file changes in a patch, making the review process more difficult and time consuming. Case in point, across the five datasets studied in section 5, we observed that 38-68% of patches are composed of more than 2 files, while in the two closed-source datasets, 25-30% of patches have more than 5 files.
Baum et al. established principles for code change ordering to help reviewers become more effective (Baum et al., 2017), while Li et al. (Li et al., 2022) and Hong et al. (Hong et al., 2022a) predict submission (patch) parts that need either review comments or a revision, yet neither of them explore or discuss the use of code diff quality as a way to prioritize the code change order. Fregnan et al.’s user study indicated that the earlier a file is shown in a patch, the more likely it is to be commented on (Fregnan et al., 2022), hence suggesting to prioritize files that need review activity to the beginning of a patch.
@@ 3.1. User study set-up
We carried out a wizard-of-oz user study (Dahlbäck et al., 1993) to compare hot-spot-based ordering of a patch’s changed files to the typical alphanumeric ordering  [^1] The study was approved by an ethical board. The certificate number is not disclosed to respect double-blind.. The wizard-of-oz design means that we used knowledge about the actual hot-spots in the oracle (obtained similarly as in Section 4.1), which is useful to evaluate the promise of hot-spot prioritization for file reordering in optimal conditions compared to the alphanumeric ordering used by most review tools (Fregnan et al., 2022). The user study was held on the premises of our industrial partner, and involved 29 developers with an average experience at the company of 6 years, knowledgeable both in developing and reviewing for the studied closed-source project, which is actively developed by more than 800 developers, with 3-4k patches per month.
@@
Each set was then associated with two patches (from the participants’ common subsystems) that none of the participants had submitted or reviewed before. We selected only patches of exactly five files, similar to Fregnan et al. (Fregnan et al., 2022). This choice is motivated by the need for patches large enough to observe the impact of ordering files while keeping the review effort manageable to attract more participants. The selected patches were not older than 2 months prior to the study, had received reviews, and needed at least one later revision or involved at least one thread of review comments.
@@ 4.1. Labeling
Our review activity prediction tasks extend existing code diff quality prediction works (Li et al., 2022; Hong et al., 2022a) in order to highlight changed files that most likely should be commented on or revised. Past approaches have used a binary labeling based on if the changed file was commented on or revised (Li et al., 2022; Hong et al., 2022a). Some works have also addressed the question of identifying useful comments, as they consider that commented changes that do not lead to revisions are not useful (Bosu et al., 2015). However, we consider such comments to still have value, as they can lead to future code submission or discussions.
Hence, in this paper, we consider similar labels (”commented” and ”revised”) as prior works, but also define a ”hot-spot” label, which indicates whether a code change has any review activity (i.e., comment, revisions, or both). This is because, while comments are meant to lead to revisions, and revisions to be caused by comments, review comments might indicate the need for revisions elsewhere in the patch or changes in future independent patches. Furthermore, we observe that 10-17% of patches across our datasets contain changes that are revised, while never commented, typically because of a review comment made on another file or outside the reviewing platform. If considering only patches that were commented on, those cases would thus be missed, hence the importance of considering hot-spot prediction. The labels’ definitions can be found in Table 3. This labeling is done at a file-level granularity, following Hong et al.’s approach (Hong et al., 2022a).
@@ 4.2. Datasets
In this paper, we use three open source datasets and two closed source datasets provided by our industrial partner (different from the dataset used in the user study). The open source datasets we consider are those of Hong et al. (Hong et al., 2022a), i.e., Qt Base (Base), Openstack Ironic (Ironic), Openstack Nova (Nova), which were used for commented and revisioned prediction tasks (Hong et al., 2022a; Tufano et al., 2021). We then used Gerrit’s API to gather extra information about the changed files, i.e., author id, reviewer ids, and full file diff (since their replication package only has the added lines), needed for our embeddings.
@@ 4.3. Data splitting
In the case of the open-source datasets, we split each dataset along its timeline into a training set validation set and test set of size 80%-10%-10% respectively, similarly to Hong et al.’s approach (Hong et al., 2022a).
@@ 4.4.1. Bag-of-Words
The BoW approach represents each input as a vector of word frequency where each column represents a word in the corpus vocabulary and the value in the vector at that column indicates the number of occurrences of the word in the input. Considering code changes as input, we have to clean the input in this context, since the source code of changes have many symbols that are not text. We follow the state-of-the-art’s approach (Hong et al., 2022a) for code pre-processing, which only considers alphanumeric terms. Numbers are also replaced by a token ”<NUMBER>”. Still following this approach, we do not perform any stemming, lowercase or lemmatization transformation as such approaches are valuable for natural languages used by humans but do not make sense in the case of code. Regarding the BoW computation itself, we use the Countvectorize function of the Scikit-Learn Python library (con, [n. d.]). The Countvectorize function is trained on the training set, then is applied to the test set for each data split. Words that appear in the test set but not in the training set will thus not be included in the bag of words. To reduce the vocabulary size, we remove words that appear only in one single changed files in the corpus as well as words that are too common (i.e., appear in more than 50% of the changed file corpus).